OpenStudy (anonymous):

graph the hyperbola -4x^2 + 9y^2 +8x +72y - 4= 0 - im not quite sure what to do since there are both x^2 and x as well as y^2 and y in this.

OpenStudy (anonymous):

complete the square

OpenStudy (anonymous):

so get y by itself? but the hyperbolas equation has both x^2 and y^2 on one side

OpenStudy (anonymous):

You have to put it in the general equation form of the hyperbola. Start by completing the square.

OpenStudy (anonymous):

k hang on im workin on it!

OpenStudy (anonymous):

Alright, first lets group the terms a little and factor some stuff:\[-4(x^{2} - 2x) +9(y^{2} +8y) = 4\] now we are going to complete the square: \[-4(x^{2} -2x +1) + 9(y^{2} +8y+16) = 4 +-4(1) + 9(16)\] this when simplified returns: \[-4(x-1)^{2}+9(y+4)^{2}=144\]

OpenStudy (anonymous):

heromiles how do you know all these helpful websites??

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

divide everything by 144: \[-(x-1)^{2}/36 + (y+4)^{2}/16 = 1\] that is standard form for a hyperbola

OpenStudy (anonymous):

ok thank you!!!!